The average number of people arriving at a fast-food drive-thru in any given 10-minute interval is 4 customers. Find the probability that at least two customers will arrive in the fast-food drive-thru in any given 15-minute interval. 


1
Expert's answer
2022-03-08T01:22:03-0500

Let X = the number of people arriving at a fast-food drive-thru in any given 15-minute interval (The interval of interest is 15 minutes)

x = 0, 1, 2, 3, …

If, the average number of people arriving at a fast-food drive-thru in any given 10-minute interval is 4 customers and there are 2/3 15 minute intervals in 10 minutes, then we have;

"\frac{1}{2/3}"(4) = 6 customers in 15 minutes, on average. So, μ = 6 for this problem. Then, the random variable X follows a Poisson distribution;

X ~ P(6)

Where;

P(x) = "\frac{\mu^x e^{-\mu}}{x!}"


We need to find P(x ≥ 2) = 1- (P(x=0) + P(x=1))


Where P(X=0) = "\frac{6^0 e^{-6}}{0!}" = 0.00248


and P(X=1) = "\frac{6^1 e^{-6}}{1!}" = 0.01487


Therefore,

 P(x ≥ 2) = 1- (0.00248+0.01487)

= 0.9827